A printer charges $1.80 for binding plus 5 cents per page for each copy of a community group's annual report.
Write the cost equation and answer the following questions.
What is the fixed cost? $1.80
what is the variable cost? .05
Let y represent the cost of the annual report and x represent the number of pages. what is the cost equation? y=1.80 + .05x
What will it cost to produce a 25 page report? $3.05
If report cost 3.30, the report is __ pages.
This last question I am having difficulty on. I have all the others right. When I try to give an answer it tells me wrong and then tells me to substitute 3.30 with the cost and solve for the length. I tried viewing examples in my program and it doesn't explain how to do this part. I am at a loss any help would be appreciated.
Thank you so much,
Dana

Answer by ramkikk66(644)

For the last part:
If the report has x pages,
the cost is 1.8 + 0.05*x (fixed binding cost + variable cost per page)
This is given to be 3.3

So 3.3 = 1.8 + 0.05*x
0.05*x = 1.5
x = 1.5/0.05 = 30
Report has 30 pages :)